# Calibration Attack: A Framework For Adversarial Attacks Targeting Calibration

### v1

@@ 4.3 Impact of Key Aspects of Attack
Table 4: Effectiveness of calibration methods and adversarial defences, evaluated with pre- and post-ECE and KS scores.
@@
Iterations. Creating attacks that are effective without requiring a large number of queries is important. Figure 2 shows the results of varying the number of iterations using the maximum miscalibration attack on the two tested models on CIFAR-100. We find that even at 100 iterations the ResNet model becomes heavily miscalibrated. In Appendix F.2, we also provide detailed comparison on the effectiveness of the maximum miscalibration attack to the base SA algorithm. We find that our attacks consistently produces higher amounts of miscalibration compared to the original SA across different number of iterations, showing how effective calibration attacks are at specifically targeting creating miscalibration, and how optimizing targeting accuracy and calibration are separate objectives and worth considering independently.
Underconfidence vs. Overconfidence. To understand which form of attack is most query efficient when the amount of change in confidence is the same, for each attack type we identify all of images in the test set that are around a given confidence level. We use the corresponding attack to made the model produce either an increase of 10% in confidence, or a decrease of 10%. We choose two base confidence levels of 80% and 90% and find all the data points within 1% of each. When an attack causes a change at or past the set threshold for the given goal probability, the attack stops and the number of queries is recorded. The results can be seen in Table 3. The consistent pattern we observe for both base confidence levels is that it takes notably fewer queries to create underconfidence than overconfidence, and the former attack is more effective at affecting the average confidence.
Qualitative Analysis. We perform qualitative analyses on the effect of calibration attacks. Appendix F.3 provides t-SNE visualization for calibration-attacked data samples. In Figure 3 below, we demonstrate the difference in the coarse localization maps using the GradCAM technique (Selvaraju et al., 2017) for images before and after calibration attack (see Appendix F.5 for details). In the first case, we confirm that all of the attacks are leading to the expected behaviour in the representation space, with the overconfidence attacks in particular being successful in pushing affected samples far from the decision boundary. With the Grad-CAM visualizations, we find that there are minimal changes to the maps after the attacks are conducted, making it difficult to identify using gradient visualization methods.
5 Results of Defending Against Calibration Attacks
In this section, we investigate 11 types of defence based on common or effective adversarial defences and calibration techniques. We compare the performance to the two methods we propose. A complete description with training details for each method is in Appendix C.2.
Baseline Defence Methods. For the calibration-focused models, we use two popular and highly effective post-calibration methods, Temperature Scaling (TS) (Guo et al., 2017) and Calibration using Splines (Gupta et al., 2021). For training-based regularization methods we use two effective models, DCA (Liang et al., 2020) and SAM (Foret et al., 2021). Regarding adversarial defence methods, we test the top-3 state-of-the-art defence models under the $l_{\infty}$ attack for CIFAR-100, using WideResNet on the RobustBench leaderboard (Croce et al., 2021). We include one of the most common defences in the form of a PGD-based adversarial training (AT) (Madry et al., 2018) baseline, which is tuned to be well calibrated on clean data, and a recent post-process defense called Adversarial Attack Against Attackers (AAA) (Chen et al., 2022), due to its effectiveness against SA.
Results and Analyses. The results for the comparison between the different methods can be seen in Table 4. We show the results of the best performing models trained in each case, and we measure query efficiency, accuracy, along with the ECE and KS error, before and after the attack. The attack model is the $l_{\infty}$ maximum miscalibration attack using the same settings described in Appendix B.
Figure 3: GradCAM visualizations on samples showing the image regions most responsible for the decisions of ResNet-50 before (top row) and after being attacked by the underconfidence and overconfidence attacks (bottom row). The left set of images shows the underconfidence attack visualizations and the right displays the overconfidence attack visualizations. Examples are picked to have large confidence differences after being attacked.
The RobustBench models compromise substantially on accuracy, and have a high level of miscalibration on clean data. They do largely avoid getting extremely miscalibrated as a result of the attacks compared to the defenceless models, except the top model on the leaderboard. Nevertheless, the high inherent miscalibration of these classifiers means they are unfavourable in situations where the model must be well calibrated.
In terms of the calibration methods, TS tends to be among the best methods at reducing calibration error prior to the attacks, but after the attacks it offers very little, if any, robustness compared to the vanilla models. The Splines method is similar in its pre-attack calibration benefits, but differs greatly in its performance post-attack. In some cases, like CIFAR-100 ResNet, it is easily the worst performing defence method. In other cases, particularly for Caltech-101 and GTSRB ViT, it is able to keep ECE at relatively reasonable values post-attack. This discrepancy shows that finding an ideal recalibration function has the potential to be a strong defence. The training-based DCA and SAM methods tend to bring few benefits after being attacked, even when they improve the calibration on clean data, the post-attack ECE and KS errors are not substantially different compared to the vanilla models.
The performance of the regular adversarial defence techniques is mixed. In terms of robustness, AAA is in most cases the technique with among the lowest post-attack ECE. Even in the best cases like Caltech-101 ResNet, the ECE tends to be at least double compared pre-attack, and in most cases we still observed multiple-fold increases. This technique is also among the poorest calibrated on clean data. Regarding AT, our approach does not compromise on accuracy and miscalibration on clean data. Likewise, it does bring some notable robustness, especially compared to the calibration methods, but it is not among the strongest.
Lastly, CAAT tends to perform more poorly in resisting the attacks than AT, likely because the training samples do not deceive the model so there is less inductive bias towards general adversarial robustness. In contrast, the performance of CS is quite strong, and it is the strongest method at maintaining low calibration error post-attack overall, even compared to the adversarial-based methods. Moreover, the technique tends to have better calibration error on clean data compared to AAA. It shows how it is key that high confidence values are retained to have decent calibration after the attacks. Altogether, despite some promising results with the defences, as a whole there are still limitations particularly with the strongest adversarial defences. The compromise of poor ECE on clean data for better calibration robustness against the attacks that we observe, as well as the general inconsistent performance means that further refinement on defences is warranted.
6 Conclusion and Outlook
We introduced and studied a new form of adversarial attack that targets the calibration of deep-learning models. Specifically, we developed four forms of calibration attacks and showed the danger of them on typical victim models, and also on models using current calibration and defence methods. We empirically showed that calibration attacks can present a major problem for safety critical systems, as they can be more difficult to detect than standard attacks, and if the calibration of the model on the data is not being closely monitored it is easy to be unaware that the model is not being attacked since the accuracy is unchanged. Also, we devised two novel defences against such calibration attacks, and empirically showed their effectiveness.
Our study here indicates that creating models that are robust against such calibration attacks through different defence methods is of tantamount importance for future work, especially with the mixed performance of various techniques we have observed, and because creating theoretically stronger and more efficient attacks is possible, particularly in terms of the overconfidence attack.
Appendix A A More Detailed Summary of Related Work
@@
The second class of methods are training-based methods, which typically add a bias during training to ensure that a model learns to become better calibrated. Often times these methods help by acting as a form of regularization that can punish high levels of overconfidence late into training. In computer vision, Mixup (Zhang et al., 2018) is a commonly used method of this type that serves as an effective regularizer by convexly combining random pairs of images and their labels and helps calibration primarily due to the use of soft, interpolated labels (Thulasidasan et al., 2019). Other methods work by adding a penalty to the loss function, like in the case of MMCE, an RKHS kernel-based measure of calibration that is added as a penalty on top of the regular loss during training so that both are optimized jointly (Kumar et al., 2018). Similarly, Tomani & Buettner (2021) create a new loss term called adversarial calibration loss that directly minimizes calibration error using adversarial samples. Given the effectiveness of many of these methods in regular testing scenarios, we desire to illustrate how well a diverse range of these methods can cope against attacks targeting model calibration and whether they possess limitations that require them to be overhauled to deal with an attack scenario.
@@
A broad range of defences against adversarial attacks have been developed, but among the most popular and effective is adversarial training (Goodfellow et al., 2015), where during training the loss is minimized over one of or both clean and generated adversarial examples. Adversarial training however greatly increases training time due to the need to fabricate adversarial examples for every batch. Gradient masking (Carlini & Wagner, 2017) is a simple defence based on obfuscating gradients so that attacks cannot make use of gradient information to create adversarial examples, although it can easily be circumvented in many cases for white-box models (Athalye et al., 2018), and black box attacks do not need gradient information in the first place. It is by and large difficult for adversarial defences to keep pace with the broad range of attacks and to be provably robust against a large number of them. Although the main topic of this work is calibration, we do focus on modelling adversarial defences and their effectiveness against our new form of attack.
Appendix B Details of Experimental Setup
Metrics. To assess the degree of calibration error caused by each attack, we use two metrics, the popular binning-based Expect Calibration Error (ECE) (Pakdaman Naeini et al., 2015), and Kolmogorov-Smirnov Calibration Error (KS error) (Gupta et al., 2021), which are formulated in detail in Section D.
@@ Appendix B Details of Experimental Setup
Models. ResNet-50 (He et al., 2016) is primary model we train and test on due to it being a standard model for image classification. Non-convolutional attention-based networks have recently attained great results on image classification tasks, so we also experiment with the popular Vision Transformer (ViT) architecture (Dosovitskiy et al., 2021). Both of these models are the versions with weights pretrained on ImageNet (Deng et al., 2009). We use the VIT_B_16 variant of ViT, and the pretraining dataset used for each model is ImageNet_1K for ResNet and ImageNet_21K for ViT, and are fine-tuned on the target datasets. Pretrained models are advantageous to study given they can increase performance over training from randomly initialized weights and is a more practical use-case. The specific details behind our training procedures and our various model hyperparameters can be seen in Section C.
Attack Settings. Regarding the SA version of the attacks, for the $l_{\infty}$ and $l_{2}$ norm attacks we use the default SA settings for $\epsilon$ and $p$, which are $\epsilon=0.05$ and $p=0.05$ for $l_{\infty}$ and $\epsilon=5.0$ and $p=0.1$ for $l_{2}$. For our primary results we run the attacks on a representative 500 test samples from the test set of each dataset. Each attack is ran for 1000 iterations, far less than the default 10,000 in Andriushchenko et al. (2020), but since there is no need to change the label, less iterations are required, bolstering the use-case and threat for this form of attacking.
@@
Appendix C Specific Training Details
C.1 General Settings
As mentioned previously, for our general attack implementation we use Square Attack, which works by using a randomized search scheme to find localized square-shaped perturbation at random positions which are sampled in such a way as to be situated approximately at the boundary of the feasible set. We still use the original sampling distributions, however we remove the initialization (initial perturbation) for each attack since it is prone to changing the predicted labels. Naturally, we use the untargeted versions of the attacks, whereby the perturbations lead to increases in the probabilites of random non-predicted classes for the underconfidence attack, since we only care about the probability of the top predicted class.
The details of the training procedure for each of the models and datasets is as follows: For CIFAR-100 and GTSRB, we use the predefined training and test sets for both but use 10% of the training data for validation purposes. For Caltech-101, which comes without predetermined splits, we use an 80:10:10 train/validation/test split. For all of the datasets, we resize all images to be 224 by 224. We also normalize all of the data based on the ImageNet channel means and standard deviations. We apply basic data augmentation during training in the form of random cropping and random horizontal flips to improve model generalizability. The hyperparameters we used for training the ResNet-50 models include: a batch size of 128, with a CosineAnnealingLR scheduler, 0.9 momentum, 5e-4 weight decay, and a stochastic gradient descent (SGD) optimizer. For ViT, the settings are the same, except we also use gradient clipping with the max norm set to 1.0. We conduct basic grid search hyperparameter tuning over a few values for the learning rate (0.1,0.01,0.005,0.001) and training duration (in terms of epochs). Generally, we found that a learning rate of 0.01 worked best for both types of models. The training times vary for each dataset and model. For the ResNet-50 models we trained for 15 epochs on CIFAR-100, 10 epochs on Caltech-101, and 7 epochs on GTSRB. Likewise for ViT, we trained for 10 epochs on CIFAR-100, 15 epochs on Caltech-101, and 5 epochs on GTSRB. The results reported in Sections 3 and 5 are shown for models on the epoch at which they attained the best accuracy on the validation set. We use the TorchVision (Paszke et al., 2019) implementation of ResNet-50, and the Huggingface Transformers (Wolf et al., 2019) implementation of ViT. All of the training occurred on 24 GB Nvidia RTX-3090 and RTX Titan GPUs. Finally, we use 15 bins to calculate the ECE.
C.2 Defense Training Settings
@@
Calibration of Neural Networks using Splines (Spline) (Gupta et al., 2021). Spine is another post-process recalibration technique that uses a recalibration function to map existing neural network confidence scores to better calibrated versions by fitting a spline function approximates the empirical cumulative distribution. It is lightweight, and often performs better than TS.
@@
Sharpness Aware Minimization (SAM) (Foret et al., 2021). SAM is a technique that improves model generalizability by simultaneously minimizing loss value and loss sharpness. It finds parameters that lie in neighbourhoods having uniformly low loss by computing the regularized ”sharpness-aware” gradient. The motivation behind using this technique as a defence is that models with parameters that lie in uniformly low loss areas may be harder to create adversarial examples, and may be more regularized. We use a neighbourhood size $\rho=0.05$. We kept the training settings the same as we described in the general settings.
@@
Calibration Attack Adversarial Training (CAAT). Our novel form of adversarial training that uses calibration attacks to generate adversarial samples rather than the regular attack algorithm. Although the general methodology is still the same as PGD-based adversarial training, the primary difference is that for each minibatch, both the underconfidence PGD calibration attack and its overconfidence version are applied to the images and the loss between the two sets of images is added. As this uses calibration attack, the labels of these images are unaffected. The settings we use for the attacks are the same as those described in B for the white-box version. Regarding the settings for each model and dataset, they are largely similar to those of regular AT, although the number of attack iterations is kept consistent at 10, even for ViT. The number of training epochs are the same as those we use for regular fine-tuning.
@@
Compression Scaling (CS). This is a novel post-process defence that does not require training and is specifically designed to maintain the regular confidence score distribution and thereby preventing extreme miscalibration while undergoing a calibration attack. Since calibration attacks does not flip the original label, For any given classifier, the strongest effect of calibration attacks will be reducing the confidence score on correctly classified “easy” samples while making the model more overconfident on difficult, misclassified samples. This creates a shift in the distribution where for a given high performing classifier the average confidence will drop dramatically while the accuracy remains high, and some misclassified samples will shift to a higher confidence level. In any case, a distribution that was originally skewed towards high confidence scores is now essentially shifted lower. Therein lies the goal of CS, to essentially shift back the distribution by scaling it such that it lies in high confidence space as before. If we assume that already low confidence correctly classified samples will be more affected by a calibration attack than one that is much higher confidence, and if we assume that incorrectly classified samples will have lower confidence then due to the relative inefficiency of the overconfidence attacks they will likely not reach extremely high confidence levels unless the attack is ran for a very large amount of iterations, then the relative ordering between many of the samples is still preserved even if the distribution is shifted, meaning the misclassified samples may still get mapped to the lower end of the confidence scale. The advantage of this method is that it largely does not incur a lot of calibration error even on clean data while being among the most effective and consistent defence methods against calibration attack. In addition, if one wants to do downstream decision making then one can still filter out the bottom $p$ percentage of images with a confidence score. For the number of bins, we mostly choose 3 or 4 as this leads to the smallest error post attack. We find the scaling factor by iterating through a large range of possible values so that the new desired confidence score for the sample is then achieved within the new confidence range.
Appendix D Calibration Metric Formulation

### v2

@@ 5 Defending Against Calibration Attacks
Table 4: Effectiveness of calibration methods and adversarial defences. The best performances of post-attack ECE (PsECE) and KS (PsKS) are marked in bold, and the second best are marked with underlines.
WideResNet
@@
Overall, from the ECE and KS scores, we can see that there are still significant limitations on recalibration and defences for calibration attacks, which invites more future research. Simple and widely used calibration models like TS are effective on clean data prior to the attacks, but they offer very little benefit post-attack. Training-based models like DCA and SAM also tend to bring few benefits after being attacked — the post-attack ECE and KS errors are not substantially different compared to the vanilla models. Lastly, we can conclude that although the defence methods from the leaderboard are generally the most adversarially resistant, their inherent high levels of miscalibration even before the attacks render them unsuitable for calibration-sensitive tasks.
6 Conclusions
We highlight and perform a comprehensive and dedicated study on calibration attacks, which aim to trap victim models into being heavily miscalibrated, hence endangering the trustworthiness of the models and any follow-up decision-making processes based on confidences. We propose four typical forms of calibration attacks and demonstrate their severity from different perspectives. We also show calibration attacks are difficult to detect compared to standard attacks. Investigation is then conducted to study the effectiveness of a wide range of adversarial defences and calibration methods, including the defences that are specifically designed for calibration attacks. From the ECE and KS scores, we can see that there are still limitations on these recalibration and defences in handling calibration attacks. We hope this paper helps attract more attention to the attacks against confidence and hence mitigate their potential harm. We provides detailed analyses to help understand the characteristics of the attacks for future work.
Appendix
Appendix A Detailed Summary of Related Work
@@
The second class of methods are training-based methods, which typically add a bias during training to ensure that a model learns to become better calibrated. Often times these methods help by acting as a form of regularization that can punish high levels of overconfidence late into training. In computer vision, Mixup (Zhang et al., 2018) is a commonly used method of this type that serves as an effective regularizer by convexly combining random pairs of images and their labels and helps calibration primarily due to the use of soft, interpolated labels (Thulasidasan et al., 2019). Other methods work by adding a penalty to the loss function, like in the case of MMCE, an RKHS kernel-based measure of calibration that is added as a penalty on top of the regular loss during training so that both are optimized jointly (Kumar et al., 2018). Similarly, Tomani & Buettner (2021) create a new loss term called adversarial calibration loss that directly minimizes calibration error using adversarial examples. Given the effectiveness of many of these methods in regular testing scenarios, we desire to illustrate how well a diverse range of these methods can cope against attacks targeting model calibration and whether they possess limitations that require them to be overhauled to deal with an attack scenario.
@@
A broad range of defences against adversarial attacks have been developed, but among the most popular and effective is adversarial training (Goodfellow et al., 2015), where during training the loss is minimized over one of or both clean and generated adversarial examples. Adversarial training however greatly increases training time due to the need to fabricate adversarial examples for every batch. Gradient masking (Carlini & Wagner, 2017) is a simple defence based on obfuscating gradients so that attacks cannot make use of gradient information to create adversarial examples, although it can easily be circumvented in many cases for white-box models (Athalye et al., 2018), and black box attacks do not need gradient information in the first place. It is by and large difficult for adversarial defences to keep pace with the broad range of attacks and to be provably robust against a large number of them. Although the main topic of this work is calibration, we do focus on modelling adversarial defences and their effectiveness against these attacks.
Appendix B Details for Maximum Miscalibration Attacks
Proof of Proposition 3.1.
Proof.
Let a classifier have non-zero accuracy. We cannot expect to reach the error of 100% since $p_{k}=0$ cannot be the case (for the top predicted class) nor can $\mathbb{P}(y_{i}=\hat{y}(\mathbf{x}_{i}))=0$ be true for all $y_{i}$. However, to achieve the highest calibration error on a set of data points in this scenario, one can first isolate the misclassified data points and if the classifier is made to output confidence scores of 100% on all of them, using the calibration attack for example, it would create a total calibration error of 100% on this set of misclassified data points. With regard to the correctly classified points, where accuracy is 100%, one can create the largest difference between the accuracy and average confidence by making the average confidence on this set as low as possible. Since confidence scores can only range from $1/K$ to 1, the largest possible difference between the average confidence score and the accuracy of 100% is $1-1/K$. Again, if $p_{k}=1/K$, and every $\hat{p}_{k}(\mathbf{x}_{i})=1/K$, while $y_{i}=\hat{y}(\mathbf{x}_{i})\forall\mathbf{x}_{i}$, then this makes the calibration error: $\mathbb{P}(y_{i}=\hat{y}(x_{i})\;|\;\hat{p}_{k}(\mathbf{x}_{i})=p_{k})-p_{k}=1-1/K$. It is not possible to create a higher level of calibration error since if $p_{k}>1/K$ on some number of the correctly classified datapoints, then $\mathbb{P}(y_{i}=\hat{y}(\mathbf{x}_{i})\;|\;\hat{p}_{k}(\mathbf{x}_{i})=p_{k})$ will still be 1, while $p_{k}>1/K$ will lead to less calibration error on that subset of datapoints. With errors on both mutually exclusive subsets of data maximized, the theoretically highest miscalibration will be created on the full data.
To derive the upper bound of the ECE value that can be achieved by MMA, if we assume $q$ is the accuracy of a $K$-way classifier $\mathcal{F}$ on the dataset $\mathcal{D}=\{\langle\mathbf{x}_{n},y_{n}\rangle\}^{N}_{n=1}$, then post successful attack all of the datapoints will fall into one of two bins representing average confidence scores of 1 and $1/K$, assuming the confidence range of each bin is 1%. The accuracy in these bins would be 0 and 1, respectively. And the proportion of data points falling into each respective bin is $1-q$ and $q$. Based on the ECE formula the maximum error would be:
$ECE_{max}$ $=\frac{n_{bin_{(100/k)\%}}}{N}|acc(bin_{(100/k)\%})-conf(bin_{(100/k)\%})|+\frac{n_{bin_{100\%}}}{N}|acc(bin_{100\%})-conf(bin_{100\%})|$ (7)
$=q*|1-1/K|+(1-q)*|0-1|$
$=1-q/K$
Appendix C Details of Experimental Setup
Metrics. To assess the degree of calibration error caused by each attack, we use two metrics, the popular binning-based Expect Calibration Error (ECE) (Pakdaman Naeini et al., 2015), and KS errors (Gupta et al., 2021), which are formulated in detail in Section E.
@@ Appendix C Details of Experimental Setup
Models. ResNet-50 (He et al., 2016) is primary model we train and test on due to it being a standard model for image classification. Non-convolutional attention-based networks have recently attained great results on image classification tasks, so we also experiment with the popular Vision Transformer (ViT) architecture (Dosovitskiy et al., 2021). Both of these models are the versions with weights pretrained on ImageNet (Deng et al., 2009). We use the VIT_B_16 variant of ViT, and the pretraining dataset used for each model is ImageNet_1K for ResNet and ImageNet_21K for ViT, and are fine-tuned on the target datasets. Pretrained models are advantageous to study given they can increase performance over training from randomly initialized weights and is a more practical use-case. The specific details behind our training procedures and our various model hyperparameters can be seen in Section D.
Attack Settings. Regarding the SA version of the attacks, for the $l_{\infty}$ and $l_{2}$ norm attacks we use the default SA settings for $\epsilon$ and $p$, which are $\epsilon=0.05$ and $p=0.05$ for $l_{\infty}$ and $\epsilon=5.0$ and $p=0.1$ for $l_{2}$. For our primary results we run the attacks on a representative 500 test cases from the test set of each dataset. Each attack is ran for 1000 iterations, far less than the default 10,000 in Andriushchenko et al. (2020), but since there is no need to change the label, less iterations are required, bolstering the use-case and threat for this form of attacking.
@@
The results in our experiment section are obtained on three runs of each model with different random seeds.
Appendix D Specific Training Details
D.1 General Settings
As mentioned previously, for our general attack implementation we use SA, which works by using a randomized search scheme to find localized square-shaped perturbation at random positions which are sampled in such a way as to be situated approximately at the boundary of the feasible set. We still use the original sampling distributions, however we remove the initialization (initial perturbation) for each attack since it is prone to changing the predicted labels. Naturally, we use the untargeted versions of the attacks, whereby the perturbations lead to increases in the probabilites of random non-predicted classes for the underconfidence attack, since we only care about the probability of the top predicted class.
The details of the training procedure for each of the models and datasets is as follows: For CIFAR-100 and GTSRB, we use the predefined training and test sets for both but use 10% of the training data for validation purposes. For Caltech-101, which comes without predetermined splits, we use an 80:10:10 train/validation/test split. For all of the datasets, we resize all images to be 224 by 224. We also normalize all of the data based on the ImageNet channel means and standard deviations. We apply basic data augmentation during training in the form of random cropping and random horizontal flips to improve model generalizability. The hyperparameters we used for training the ResNet-50 models include: a batch size of 128, with a CosineAnnealingLR scheduler, 0.9 momentum, 5e-4 weight decay, and a stochastic gradient descent (SGD) optimizer. For ViT, the settings are the same, except we also use gradient clipping with the max norm set to 1.0. We conduct basic grid search hyperparameter tuning over a few values for the learning rate (0.1,0.01,0.005,0.001) and training duration (in terms of epochs). Generally, we found that a learning rate of 0.01 worked best for both types of models. The training times vary for each dataset and model. For the ResNet-50 models we trained for 15 epochs on CIFAR-100, 10 epochs on Caltech-101, and 7 epochs on GTSRB. Likewise for ViT, we trained for 10 epochs on CIFAR-100, 15 epochs on Caltech-101, and 5 epochs on GTSRB. The results reported in Sections 3 and 5 are shown for models on the epoch at which they attained the best accuracy on the validation set. All of the training occurred on 24 GB Nvidia RTX-3090 and RTX Titan GPUs. Finally, we use 15 bins to calculate the ECE.
D.2 Defence Training Settings
@@
Multi-domain Temperature Scaling (MD-TS) (Yu et al., 2022) MD-TS is a method based on TS but is designed to be more robust in situations when data comes from multiple domains, as in this case with images corrupted using different types of calibration attacks. It modifies the original TS method by first finding the ideal temperature across each domain, then training a linear regression classifier using the feature embeddings of each datapoint and the corresponding ideal temperatures based on the respective domain, yielding a classifier that can dynamically calculate an ideal temperature for each instance at test time. We modify this domain for our task of defence by creating three different domains for the base images and their underconfidence attacked and overconfidence attacked counterparts. We select 500 validation instances and find the temperature for each domain and conduct the rest of the method as in its original incarnation. The feature embeddings are as before, using the penultimate layer outputs of model before the classification layer. We experiment with converting the feature embeddings to Fourier domain using the fast Fourier transform before feeding them to the classifier, similar to the principle behind the detection method in Harder et al. (2021) to make it easier to identify adversarially attacked datapoints, though we find that the conversion brings little benefit over the base variation of the method.
Calibration of Neural Networks using Splines (Spline) (Gupta et al., 2021). Spline is another post-process recalibration technique that uses a recalibration function to map existing neural network confidence scores to better calibrated versions by fitting a spline function approximates the empirical cumulative distribution. It is lightweight, and often performs better than TS.
@@
Sharpness Aware Minimization (SAM) (Foret et al., 2021). SAM is a technique that improves model generalizability by simultaneously minimizing loss value and loss sharpness. It finds parameters that lie in neighbourhoods having uniformly low loss by computing the regularized "sharpness-aware" gradient. The motivation behind using this technique as a defence is that models with parameters that lie in uniformly low loss areas may be harder to create adversarial examples, and may be more regularized. We use a neighbourhood size $\rho=0.05$. We kept the training settings the same as we described in the general settings.
@@
Calibration Attack Adversarial Training (CAAT). Our novel form of adversarial training that uses calibration attacks to generate adversarial examples rather than the regular attack algorithm. Although the general methodology is still the same as PGD-based adversarial training, the primary difference is that for each minibatch, both the underconfidence PGD calibration attack and its overconfidence version are applied to the images and the loss between the two sets of images is added. As this uses calibration attack, the labels of these images are unaffected. The settings we use for the attacks are the same as those described in C for the white-box version. Regarding the settings for each model and dataset, they are largely similar to those of regular AT, although the number of attack iterations is kept consistent at 10, even for ViT. The number of training epochs are the same as those we use for regular fine-tuning.
@@
Compression Scaling (CS). This is a novel post-process defence that does not require training and is specifically designed to maintain the regular confidence score distribution and thereby preventing extreme miscalibration while undergoing a calibration attack. Since calibration attacks does not flip the original label, for any given classifier, the strongest effect of calibration attacks will be reducing the confidence score on correctly classified “easy” datapoints while making the model more overconfident on difficult, misclassified datapoints. This creates a shift in the distribution where for a given high performing classifier the average confidence will drop dramatically while the accuracy remains high, and some misclassified datapoints will shift to a higher confidence level. In any case, a distribution that was originally skewed towards high confidence scores is now essentially shifted lower. Therein lies the goal of CS, to essentially shift back the distribution by scaling it such that it lies in high confidence space as before. If we assume that already low confidence correctly classified datapoints will be more affected by a calibration attack than one that is much higher confidence, and if we assume that incorrectly classified datapoints will have lower confidence then due to the relative inefficiency of the overconfidence attacks they will likely not reach extremely high confidence levels unless the attack is ran for a very large amount of iterations, then the relative ordering between many of the datapoints is still preserved even if the distribution is shifted, meaning the misclassified datapoints may still get mapped to the lower end of the confidence scale. The advantage of this method is that it largely does not incur a lot of calibration error even on clean data while being among the most effective and consistent defence methods against calibration attack. In addition, if one wants to do downstream decision making then one can still filter out the bottom $p$ percentage of images with a confidence score. For the number of bins, we mostly choose 3 or 4 as this leads to the smallest error post attack. We find the scaling factor by iterating through a large range of possible values so that the new desired confidence score for the datapoint is then achieved within the new confidence range.
Binning Details. In our two defence algorithms, the range of possible confidence scores are first split into equally sized bins. In our case, we divide confidence scores into 15 bins and chose the top 3 (or 4) highest confidence bins as the compressed bins as mentioned above.
Appendix E Calibration Metric Formulation